A 590 g squirrel with a surface area of 905 cm2 falls from a 4.4-m tree to the ground. Estimate its terminal velocity. (Use the drag coefficient for a horizontal skydiver. Assume that the squirrel can be approximated as a rectanglar prism with cross-sectional area of width 11.4 cm and length 22.8 cm. Note, the squirrel may not reach terminal velocity by the time it hits the ground. Give the squirrel's terminal velocity, not it's velocity as it hits the ground.
Answer #1

Terminal velocity is reached when weight is equal to drag force

mg = 1/2v2ACd

where is density of air, v is velocity, A is cross section area , Cd is drag coefficient

v = sqrt ( 2mg / ACd)

drag coefficient for horizontal diver = 1

v = sqrt ( 2 * 590e-3 * 9.8 / 1.21 * 11.4e-2 * 22.8e-2 * 1)

v = sqrt ( 11.564 / 0.03145)

v = 19.17 m/s